OpenStudy (anonymous):

how do I solve the inequality 3a-4 less than or equal to 5 I know the first step would be add 4 to each side

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

Ok!, try to add +4 to both sides of your inequality, please

OpenStudy (anonymous):

\[3a-4\le5 Afteradd +4 \to both side then: devide \to 3 then we have: a \le \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

sorry \[a \le \frac{ 9 }{ }\]

OpenStudy (anonymous):

I did and got 3a greater then or equal to 9

OpenStudy (anonymous):

\[a \le \frac{ 9 }{ 3}\]

OpenStudy (michele_laino):

@olivia.smiles here is the procedure: \[3a-4+4\le 5+4\]

OpenStudy (anonymous):

@Michele_Laino \[3a \le 9\]

OpenStudy (michele_laino):

well done!

OpenStudy (michele_laino):

now divide by 3, your last inequality, please

OpenStudy (anonymous):

a is greater then or equal to 3

OpenStudy (michele_laino):

are you sure? I think is: \[\frac{ 3a }{ 3 }\le \frac{ 9 }{ 3 }\], then?

OpenStudy (solomonzelman):

you put the wrong sign, olivia.smiles, you said that "a" is GREATER than or equal to 3, but it is LESS than or equal to 3.

OpenStudy (michele_laino):

@SolomonZelman from the text I read a is less or equal, so the sigh is \[\le \]

OpenStudy (solomonzelman):

I said that it is less than or equal to. Because Olivia said that it is greater or equal.

OpenStudy (michele_laino):

@SolomonZelman Sorry I did not understand!

OpenStudy (solomonzelman):

It's totally fine:)

OpenStudy (michele_laino):

@SolomonZelman thanks!

OpenStudy (solomonzelman):

nothing to thank me for, but you are always welcome!

OpenStudy (michele_laino):

@SolomonZelman thanks again!

OpenStudy (solomonzelman):

:)